Radon levels in Egyptian tombs 509101805807800 4000 390 12100 3400 1300 11900 1100Construct a 90% confidence interval for \eta
Answer #1

Answer)

As the population s.d is unknown here we will use t distribution to estimate the interval

First we need to find the mean and s.d of the given data

Mean = 3642.5

S.d = 4486.929

N = 12

Degrees of freedom is = n-1 = 11

For 11 df and 90% confidence level

Critical value t from t table is = 1.8

Margin of error (MOE) = t*s.d/√n = 1.8*4486.929/√12 = 2326.146

Confidence interval is given by

(Mean - Moe, Mean + MOE)

(1316.354, 5968.646)